OpenStudy (anonymous):

Simplify: 2x - 2x^2 / x^3 - 2x^2 + x

OpenStudy (saifoo.khan):

Factorize.

OpenStudy (anonymous):

Can you solve the problem please? I factored it but it looks...wrong.

OpenStudy (saifoo.khan):

\[\frac{2x(1- x)}{x(x-1)(x-1)}=\frac{-2x(-1+ x)}{x(x-1)(x-1)}=\frac{-2x\cancel{(-1+ x)}}{x(x-1)\cancel{(x-1)}}\]

OpenStudy (saifoo.khan):

\[\frac{-2x}{x(x-1)}\]

hero (hero):

Wow. I had no clue a cancel function existed

OpenStudy (anonymous):

Thanks for the help! I was on the right track after all...

OpenStudy (saifoo.khan):

you welcome.

OpenStudy (anonymous):

Although I do have a question. In the second step of the, why did you make 2x negative while in the parenthesis you changed the sign of 1 and x? Is it necessary to solve the problem?

OpenStudy (saifoo.khan):

otherwise you cannot cancel them.